# On the regularity of axisymmetric, swirl-free solutions of the Euler equation in four and higher dimensions

### v3

@@ 6 Anti-parallel vortex tubes
$\int_{0}^{\infty}\int_{0}^{\infty}r^{d-2}z\omega u_{r}\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z$ $=\int_{0}^{\infty}\int_{0}^{\infty}r^{d-2}zu_{r}(\partial_{r}u_{z}-\partial_{z}u_{r})\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z$ (6.80)
$=\int_{0}^{\infty}\int_{0}^{\infty}\left(-zu_{z}\partial_{r}\left(r^{d-2}u_{r}\right)-\frac{1}{2}r^{d-2}z\partial_{z}u_{r}^{2}\right)\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z$ (6.81)
$=\int_{0}^{\infty}\int_{0}^{\infty}\left(zu_{z}\partial_{z}\left(r^{d-2}u_{z}\right)+\frac{1}{2}r^{d-2}u_{r}^{2}\right)\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z$ (6.82)
$=\frac{1}{2}\int_{0}^{\infty}\int_{0}^{\infty}\left(-r^{d-2}u_{z}^{2}+r^{d-2}u_{r}^{2}\right)\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z.$ (6.83)
@@
$\int_{0}^{\infty}\int_{0}^{\infty}r^{d-1}\omega u_{z}\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z$ $=\int_{0}^{\infty}\int_{0}^{\infty}r^{d-1}(\partial_{r}u_{z}-\partial_{z}u_{r})u_{z}\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z$ (6.84)
$=-\frac{d-1}{2}\int_{0}^{\infty}\int_{0}^{\infty}r^{d-2}u_{z}^{2}\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z+\int_{0}^{\infty}\int_{0}^{\infty}r^{d-1}u_{r}\partial_{z}u_{z}\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z.$ (6.85)
@@ Appendix A A stream function approach to the Biot-Savart Law
$u_{z}(r,z)=\frac{1}{2\pi}\iint_{\Pi}\int_{0}^{\pi}(\sin\theta)^{k-1}\cos\theta\left(\frac{k(r-\bar{r}\cos\theta)}{X^{\frac{k}{2}+1}}-\frac{k}{rX^{\frac{k}{2}}}\right)d\theta\,\omega(\bar{r},\bar{z})\bar{r}^{k}\,d\bar{r}\,d\bar{z}=I_{1}+I_{2}.$ (A.13)
@@
$\begin{split}u_{z}(r,z)&=I_{1}+I_{2}\\ &=\frac{k}{2\pi}\iint_{\Pi}\int_{0}^{\pi}\frac{(\sin\theta)^{k-1}}{X^{\frac{d}{2}}}\bigg[(\cos\theta)(r-\bar{r}\cos\theta)-(\sin\theta)^{2}\bar{r}\bigg]\,d\theta\,\omega(\bar{r},\bar{z})\bar{r}^{k}\,d\bar{r}\,d\bar{z}\\ &=\frac{k}{2\pi}\iint_{\Pi}\int_{0}^{\pi}\frac{(\sin\theta)^{k-1}}{X^{\frac{d}{2}}}(r\cos\theta-\bar{r})\,d\theta\,\omega(\bar{r},\bar{z})\bar{r}^{k}\,d\bar{r}\,d\bar{z}.\end{split}$ (A.15)
@@ Acknowledgements
Both authors were partially supported by NSERC under the grants RGPIN-2023-04534 and RGPIN-2018-04137. EM was also supported by the Pacific Institute for the Mathematical Sciences as a PIMS postdoctoral fellow at the University of British Columbia. This PIMS postdoctoral fellowship was supported by NSERC under grant no. 568577-2022. The authors would like to thank Stephen Gustafson for his helpful advice on this problem. The authors would also like to thank the referees for their thorough reading of the manuscript and helpful suggestions.

### v4

@@ 6 Anti-parallel vortex tubes
$\int_{0}^{\infty}\int_{0}^{\infty}r^{d-2}z\omega u_{r}\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z$ $=\int_{0}^{\infty}\int_{0}^{\infty}r^{d-2}zu_{r}(\partial_{r}u_{z}-\partial_{z}u_{r})\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z$
$=\int_{0}^{\infty}\int_{0}^{\infty}\left(-zu_{z}\partial_{r}\left(r^{d-2}u_{r}\right)-\frac{1}{2}r^{d-2}z\partial_{z}u_{r}^{2}\right)\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z$
$=\int_{0}^{\infty}\int_{0}^{\infty}\left(zu_{z}\partial_{z}\left(r^{d-2}u_{z}\right)+\frac{1}{2}r^{d-2}u_{r}^{2}\right)\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z$
$=\frac{1}{2}\int_{0}^{\infty}\int_{0}^{\infty}\left(-r^{d-2}u_{z}^{2}+r^{d-2}u_{r}^{2}\right)\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z.$
@@
$\int_{0}^{\infty}\int_{0}^{\infty}r^{d-1}\omega u_{z}\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z$ $=\int_{0}^{\infty}\int_{0}^{\infty}r^{d-1}(\partial_{r}u_{z}-\partial_{z}u_{r})u_{z}\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z$
$=-\frac{d-1}{2}\int_{0}^{\infty}\int_{0}^{\infty}r^{d-2}u_{z}^{2}\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z+\int_{0}^{\infty}\int_{0}^{\infty}r^{d-1}u_{r}\partial_{z}u_{z}\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z.$
@@ Appendix A A stream function approach to the Biot-Savart Law
$u_{z}(r,z)$ $=\frac{1}{2\pi}\iint_{\Pi}\int_{0}^{\pi}(\sin\theta)^{k-1}\cos\theta\left(\frac{k(r-\bar{r}\cos\theta)}{X^{\frac{k}{2}+1}}-\frac{k}{rX^{\frac{k}{2}}}\right)d\theta\,\omega(\bar{r},\bar{z})\bar{r}^{k}\,d\bar{r}\,d\bar{z}$
$=I_{1}+I_{2}.$
@@
$u_{z}(r,z)$ $=I_{1}+I_{2}$
$=\frac{k}{2\pi}\iint_{\Pi}\int_{0}^{\pi}\frac{(\sin\theta)^{k-1}}{X^{\frac{d}{2}}}\bigg[(\cos\theta)(r-\bar{r}\cos\theta)-(\sin\theta)^{2}\bar{r}\bigg]\,d\theta\,\omega(\bar{r},\bar{z})\bar{r}^{k}\,d\bar{r}\,d\bar{z}$
$=\frac{k}{2\pi}\iint_{\Pi}\int_{0}^{\pi}\frac{(\sin\theta)^{k-1}}{X^{\frac{d}{2}}}(r\cos\theta-\bar{r})\,d\theta\,\omega(\bar{r},\bar{z})\bar{r}^{k}\,d\bar{r}\,d\bar{z}.$
@@ Acknowledgements
The authors would like to thank Stephen Gustafson for his helpful advice on this problem. The authors would also like to thank the referees for their thorough reading of the manuscript and helpful suggestions.
Both authors were partially supported by NSERC under the grants RGPIN-2023-04534 and RGPIN-2018-04137. EM was also supported by the Pacific Institute for the Mathematical Sciences as a PIMS postdoctoral fellow at the University of British Columbia. This PIMS postdoctoral fellowship was supported by NSERC under grant no. 568577-2022.